OpenStudy (anonymous):

If each portrait is offered in one of 5 different sizes, one of 3 different colors and one of 8 different borders, how many different portrait combinations are there?

OpenStudy (anonymous):

why do you delete your questions ?

jimthompson5910 (jim_thompson5910):

There are 5*3*8 = 15*8 = 120 different combinations